gelphielvr:

(algebra 1) need help turning this function into standard form. function = h(x)= a(x+5)-3 The vertex is (-5,-3)

gelphielvr:

if you need more info I can post the graph but I honestly have no clue on what to do

Hero:

Pretty sure your function is supposed to be \(h(x)=a(x+5)^2-3\)

gelphielvr:

oh my bad, I forgot to put the squared 2.

Hero:

Yes and then just replace \(h(x)\) with \(y\)

gelphielvr:

@hero wrote:
Yes and then just replace \(h(x)\) with \(y\)
so it'd be y = a(x=5)-3 ? do you know what I have to do with the a

bubbly1belle:

You want to try to simplify it as much as possible. use the exponent and distribute it. You next want to distribute the 'a'. then try to order the equation so the variables squared in front and the regular numbers at the end.

bubbly1belle:

ex: y= x^2+x+2

Hero:

\(y=a(x+5)^2-3\) Plug the vertex \((x,y=(-5,-3)\) in then solve for a

bubbly1belle:

is the function being solved or just put into standard form?

Hero:

@bubbly1belle wrote:
is the function being solved or just put into standard form?
You have to find \(a\) first

gelphielvr:

@hero wrote:
@bubbly1belle wrote:
is the function being solved or just put into standard form?
You have to find \(a\) first
alright, thanks for helping a gender "confused" person